OpenStudy (anonymous):

List the discontinuities for the function f(x) = cot(5x/3 ).

OpenStudy (anonymous):

\[\cot\frac{ 5x }{ 3 }\]

OpenStudy (cwrw238):

what is the value when x = 0?

OpenStudy (anonymous):

THESE are the answeres: a. There are no discontinuities. b. n(pi/2), where n is an integer c. n(pi/3), where n is an integer d. n(three(pi)/5), where n is an integer

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

@l_o_w_key @cwrw238

myininaya (myininaya):

cot(x)=cos(x)/sin(x)... you need to basically solve \[\sin(\frac{5x}{3})=0\]

OpenStudy (anonymous):

what happened to cos?

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

you can't divide by 0

myininaya (myininaya):

you have sin(5x/3) on bottom

myininaya (myininaya):

you cannot have sin(5x/3)=0

OpenStudy (anonymous):

i still dont get it

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

\[\cot(\frac{5x}{3})=\frac{\cos(\frac{5x}{3})}{\sin(\frac{5x}{3})}\] agreed?

OpenStudy (anonymous):

ya

OpenStudy (solomonzelman):

no

myininaya (myininaya):

a fraction is discontinuous when its bottom is 0

myininaya (myininaya):

\[\sin(\frac{5x}{3})=0 \text{ will give discontinuites }\]

OpenStudy (solomonzelman):

oh, sorry

OpenStudy (solomonzelman):

it's cotangent, I thought it was cosine

OpenStudy (solomonzelman):

I should pay more attention to the questions.

OpenStudy (anonymous):

okay so how do i solve the problem to get the answer?

myininaya (myininaya):

let u=5x/3 do you know how to solve sin(u)=0

myininaya (myininaya):

when you are done solve sin(u)=0 for u then you can finish up by replacing u with 5x/3 then solve for x

OpenStudy (anonymous):

i dont know how to do tht

OpenStudy (anonymous):

how do u do tht

myininaya (myininaya):

go to your unit circle and find for what angles is sine (the y-coordinate ) 0

OpenStudy (anonymous):

squareroot 3/2

myininaya (myininaya):

that's an angle?

myininaya (myininaya):

I think you just randomly chose a value within one of the ordered pairs

myininaya (myininaya):

I'm asking you to find angle(s) on the unit circle for which the y-coordinate is 0 (0 means zero)

OpenStudy (anonymous):

120 and 300

myininaya (myininaya):

sin(120) is not zero sin(300) is also not zero

OpenStudy (anonymous):

ohh okay so its 180 and 360 .

OpenStudy (anonymous):

???

myininaya (myininaya):

yes and you can also so 0 right?

myininaya (myininaya):

0,180,360,... or 180*n where n is an integer

myininaya (myininaya):

I think they want it in radians so convert that to radians

myininaya (myininaya):

\[\frac{5x}{3}=\pi n\]

myininaya (myininaya):

finish solving for x and you are done

OpenStudy (anonymous):

so its pi/2?

myininaya (myininaya):

trying to think how you got that after solving for x

myininaya (myininaya):

to undo the division by 3 multiply by 3 on both sides to undo the multiplication by 5 you divide both sides by 5

myininaya (myininaya):